OpenStudy (anonymous):

Last week Brandon and Marcee paid a total of $16 to go to the movies. This week three of their friends are joining them. How much can they expect to pay altogether?

OpenStudy (anonymous):

$40.00 Dollars :)

OpenStudy (anonymous):

THankkk you so much:) can you help me with more?

OpenStudy (anonymous):

Ya sure Anytime

OpenStudy (anonymous):

thats my next question

OpenStudy (anonymous):

I would say 85 degrees

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Sorry but im not so sure about this one its not A that narrows it down a little sorry man thats all i got

OpenStudy (anonymous):

no worries!!!! :)

OpenStudy (anonymous):

not 30/32 not 5/6 I would say 4/6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I WOULD SAY THE BOTTOM ONE

OpenStudy (anonymous):

okey dokey

OpenStudy (anonymous):

hey I got to get going but u can ask a guy his name is Blank he's really good ;)

OpenStudy (anonymous):

and he's on right now

OpenStudy (anonymous):

how do i ask him

OpenStudy (anonymous):

@Blank  hi can you help me

OpenStudy (anonymous):

through chat you can message him through the messaging system its on the left hand corner its the envelope. :)

OpenStudy (blank ):

B. 12/7 = 18/a

OpenStudy (anonymous):

?

OpenStudy (blank ):

The answer is D. s/4 = 9.8/7